Using a penny and a six-sided number cube:
88. P(heads)
89.P(even)
90.P(tails and 3)

Answer by stanbon(75887)

Using a penny and a six-sided number cube:
88. P(heads) = 1/2
----
89.P(even) = (3/6) = 1/2
--------
90.P(tails and 3) = (1/2)*(1/6) = 1/12
============================================
Cheers,
Stan H.